# Diffusion approximation of the stationary distribution of a two-level single server queue

### v5

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under the scaling in heavy traffic which corresponds to a diffusion approximation. It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [2, 3, 10, 11]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distributions.
Keywords: single server queue, level dependent arrival and service, stationary distribution, heavy traffic, diffusion approximation, Palm distribution, process limit, stochastic integral equation.
11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan.
@@ 1 Introduction
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. It may be considered to approximate it by a quasi-birth-and-death process, then it is not hard to numerically compute its performance. However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential. Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue (e.g., see Kingman [7]). Furthermore, Miyazawa [11] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see Harrison [5]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [2, 3, 10, 11]. This verification, Theorem 3.1, requires either some limitation on system parameters or for the workload brought by a customer to be exponentially distributed. The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
@@
2 Framework for the 2-level GI/G/1 queues
@@ 3 A sequence of the 2-level GI/G/1 queues
We consider the scaling limit of the stationary distribution of the 2-level $GI/G/1$ queue in heavy traffic. To this end, we introduce a sequence of this queueing models indexed by $r\in(0,1]$. Since we consider their countable sequence, one may consider $r=1/m$ for positive integer $m$, but we keep $r$ just for convenience.
@@ 3.1 Assumptions
For the $r$-th model, let $\ell^{(r)}_{0}$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$. We refer to this number as the queue length. Let $t^{(r)}_{e,n}$ be the $n$-th arrival time of a customer, and let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customer. We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$. Thus, the inter-arrival times are independent, but they may have two different distributions depending on the queue length. Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are random variables which may have different distributions. As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers. We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variable. We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$. Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{0}$ and otherwise, respectively. Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
@@
    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$, and $\ell^{(r)}_{0}\geq 2$.
@@
Here, we write $f(r)=o(g(r))$ as $r\downarrow 0$ for positive valued functions $f,g$ of $r$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$. Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$. Note that (3.a) implies that
@@
We take countably many $r$’s for a limiting operation. For such a sequence, there exists a stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator $\Theta_{\bullet}$ on which $X^{(r)}(\cdot)$ is a stationary process. In what follows, $\mathbb{P}$ is always this probability measure. By the condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution. Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$. For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$. For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting distribution agrees with that of $T^{(r)}_{e,i}$.
@@
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}(n)\wedge T_{e,2}(n)]<\infty$. Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)]<\infty$. ∎
@@ 3.2 Main results
We refer to this $\nu^{(r)}$ as a scaled stationary distribution. We decompose $\nu^{(r)}$ into measures $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
@@
Conversely, if (3.12) holds, then we have the weak convergence (3.11) from (i) because $\widetilde{A}_{i}>0$. Thus, (3.12) is equivalent to (3.11). We conjecture that (3.11) holds without any condition in (ii). There are two supports for this conjecture. One is from a process limit for the 2-level $GI/G/1$ queue in heavy traffic. Another is from the intuitive verification of (3.12). They are detailed in Sections 5.1 and 5.2, repectively.
@@
Let consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its direct proof in Appendix A.2, which is helpful to intuitively convince (ii) of Theorem 3.1. This proof is straightforward, but one must be careful because $L^{(r)}(\cdot)$ itself is not a continuous-time Markov chain even under the exponential assumptions (see Appendix A.2 for its reason).
@@ 4.1 Palm distribution and BAR
[…]
@@
We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [11]. For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$. Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded functions from $S$ to $\mathbb{R}_{+}$ which continuously and partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right. Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
@@
In what follows, we often use a difference operator $\Delta$, which is defined as $\Delta g(t)=g(t)-g(t-)$ for function $g$ from $\mathbb{R}_{+}$ to $\mathbb{R}$ which is right continuous with the left-limit. Define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
@@
For convenience, we define another difference operators $\Delta_{e}$ and $\Delta_{s}$ as
@@
where the summation is well defined because $\Delta_{e}f(X^{(r)}(u))$ and $\Delta_{s}f(X^{(r)}(u))$ vanish except for finitely many $u\in(0,t]$. Taking the expectation of the above equation by $\mathbb{P}$, we have a stationary equation:
@@
    $X^{(r)}(0-)$ is not defined for $X^{(r)}(\cdot)=\{X^{(r)}(t);t\geq 0\}$, but well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$. Thus, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined.
@@
    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00. This means that we take a so called arrival first framework. As for the intermediate state, see [11] for its details. We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$. This prevents to explicitly refer to the intermediate state $X^{(r)}_{1}(0)$. Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
@@
Apply this to (4.1) yields

### v6

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of the queue length in this two-level queue under scaling in heavy traffic. This limit corresponds to its diffusion approximation. It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [2, 3, 10, 12]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distribution under scaling.
Keywords: single server queue; level dependent arrival and service; stationary distribution; heavy traffic; diffusion approximation; Palm distribution; stochastic integral equation; multi-level SRBM.
@@ 1 Introduction
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured. Because of these facts, an approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see Kingman [7]). Furthermore, Miyazawa [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see Harrison [5]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [2, 3, 10, 12]. This verification, Theorem 3.1, requires either some limitation on system parameters or for the workload brought by a customer to be exponentially distributed. The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
@@
2 Framework for the 2-level GI/G/1 queue
@@ 3 A sequence of the 2-level GI/G/1 queues
We consider the scaling limit of the stationary distribution of the 2-level $GI/G/1$ queue in heavy traffic. To this end, we introduce a sequence of this queueing models indexed by $r\in(0,1]$. Since this sequence can be countable, one may take $r=1/\sqrt{n}$ for positive integer $n$ so that the scaling is identical with the standard diffusion approximation, but we keep $r$ just for convenience.
@@ 3.1 Assumptions
For the $r$-th model, let $\ell^{(r)}_{0}\geq 2$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$. We refer to this number as the queue length. Let $t^{(r)}_{e,n}$ be the $n$-th arrival time of a customer, and let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customer. We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$. Thus, the inter-arrival times are independent, but they may have two different distributions depending on the queue length. Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are independent random variables which may have different distributions. As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers. We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variable. We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$. Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{0}$ and otherwise, respectively. Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
@@
    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$.
@@
Here, for positive valued functions $f,g$ of $r$, we write $f(r)=o(g(r))$ as $r\downarrow 0$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$. Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$. Note that (3.a) implies that
@@
We take countably many $r$’s for a limiting operation. For such a sequence, there exists a stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator $\Theta_{\bullet}$ on which $X^{(r)}(\cdot)$ is a stationary process. In what follows, $\mathbb{P}$ is always this probability measure. By the condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution. Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$. For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$. For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting times have the same distribution as that of $T^{(r)}_{e,i}$.
@@
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$. Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}]<\infty$. ∎
@@ 3.2 Main results
We refer to this $\nu^{(r)}$ as a scaled stationary distribution. We decompose $\nu^{(r)}$ using finite measures $\nu^{(r)}_{i}$ and conditional probability distributions $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$, which are defined as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
@@
Conversely, if (3.12) holds, then we have the weak convergence (3.11) from (i) because $\widetilde{A}_{i}>0$. Thus, (3.12) is equivalent to (3.11). We conjecture that (3.11) holds without any condition in (ii). This is supported by the intuitive verification of (3.12), which is detailed in Section 5.2.
@@
Let us consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its direct proof in Appendix A.2, which is helpful to convince (ii) of Theorem 3.1. This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of more than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1). This Markov chain will be detailed in Appendix A.2.
@@ 4.1 Palm distribution and BAR
Then, for each $r\in(0,1]$, $X^{(r)}(\cdot)$ and the increments of $N^{(r)}_{e}$ and $N^{(r)}_{s}$ are jointly stationary. Since $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite by Lemma 3.1, we can define probability distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ on $(\Omega,{\mathcal{F}})$ by
@@
We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [12]. For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$. Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded and continuous functions from $S$ to $\mathbb{R}_{+}$ which are partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right. Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
@@
In what follows, we often use a difference operator $\Delta$, which is defined as $\Delta g(t)=g(t)-g(t-)$ for function $g$ from $\mathbb{R}_{+}$ to $\mathbb{R}$ which is right continuous with the left-limit. For $t\geq 0$, define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
@@
In these definitions, $X^{(r)}(0-)$ is required, but formally not defined. However, this does not cause any problem. Please see (i) of Remark 4.1 for its details. For convenience, we define another difference operators $\Delta_{e}$ and $\Delta_{s}$ as
@@
where the summation is well defined because $\Delta_{e}f(X^{(r)}(u))$ and $\Delta_{s}f(X^{(r)}(u))$ vanish except for finitely many $u\in(0,t]$. Take the expectation of the above equation for $t=1$ by $\mathbb{P}$, then the stationarity of $X(\cdot)$ and the definition of the Palm distribution yield
[…]
$\mathbb{E}\left[\sum_{u\in(0,1]}\Delta_{v}f(X^{(r)}(u))]\right]=\alpha^{(r)}_{v}\mathbb{E}^{(r)}_{v}\left[\Delta_{v}f(X^{(r)}(0))\right],\qquad v=e,s.$
Hence, we have a stationary equation:
@@
    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distribution $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$. So, $X^{(r)}(0-)$ can be considered as a random variable generated by $\mathbb{P}^{(r)}_{e}$. However, one may be uncomfortable for this interpretation. In this case, please consider the process $X(t)$ to start at time $-1$ subject to the stationary distribution, then the distribution of $X(\cdot)=\{X(t);t\geq 0\}$ is unchanged because $X(\cdot)$ is a stationary process. In either way, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions.
@@
    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00. This means that we take a so called arrival first framework. As for the intermediate state, see [12] for its details. We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$. This prevents to directly refer to the intermediate state $X^{(r)}_{1}(0)$ in our arguments. Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
@@
Applying this to (4.1) yields